# ControlText: Unlocking Controllable Fonts in Multilingual Text Rendering without Font Annotations

### v1

@@ Abstract
This work demonstrates that diffusion models can achieve font-controllable multilingual text rendering using just raw images without font label annotations. Visual text rendering remains a significant challenge. While recent methods condition diffusion on glyphs, it is impossible to retrieve exact font annotations from large-scale, real-world datasets, which prevents user-specified font control. To address this, we propose a data-driven solution that integrates the conditional diffusion model with a text segmentation model, utilizing segmentation masks to capture and represent fonts in pixel space in a self-supervised manner, thereby eliminating the need for any ground-truth labels and enabling users to customize text rendering with any multilingual font of their choice. The experiment provides a proof of concept of our algorithm in zero-shot text and font editing across diverse fonts and languages, providing valuable insights for the community and industry toward achieving generalized visual text rendering.
@@
Bowen Jiang1thanks: Co-first authors. This is preliminary work and code will be released at github.com/bowen-upenn/ControlText. , Yuan Yuan [^1] footnotemark: 1 , Xinyi Bai2, Zhuoqun Hao1, Alyson Yin3, Yaojie Hu1, Wenyu Liao1, Lyle Ungar1, Camillo J. Taylor1 University of Pennsylvania1 Cornell University2 University of California, Irvine3 Philadelphia, PA 19104 Ithaca, NY 14850 Irvine, CA 92697 bwjiang, yyuan86@seas.upenn.edu, ungar@cis.upenn.edu, cjtaylor@seas.upenn.edu
@@ 1 Introduction
Diffusion models have become the dominant paradigm in image generation Ho et al. (2020); Saharia et al. (2022); Zhang et al. (2023b); Rombach et al. (2022); Betker et al. (2023); Ramesh et al. (2022); Esser et al. (2024), because of their iterative denoising process that allows fine-grained image synthesis. While these models effectively capture data distributions of photorealistic or artistic images, they still fall short in generating high-fidelity text. Rendering text in images is inherently more challenging as it requires precise knowledge of the geometric alignment among strokes, the arrangement of letters as words, the legibility across varying fonts, sizes, and styles, and the integration of text into visual backgrounds. At the same time, humans are more sensitive to minor errors in text, such as a missing character or an incorrectly shaped letter, compared to natural elements in a visual scene that allow for a much higher degree of variation.
@@
  • There are numerous fonts available in the open world, but many fonts with different names are very similar which confounds evaluation.
@@
  • State-of-the-art image generation models are either closed-sourced Betker et al. (2023); Midjourney or users still cannot edit texts or fonts without altering other parts of the image.
@@
We also provide the community a comprehensive dataset for font-aware glyph controls collected from diverse real-world images. We further propose a quantitative evaluation metrics for handling fuzzy fonts in the open world. Experimental results demonstrate that our method, ControlText, as a text and font editing model, facilitates a human-in-the-loop process to generate multilingual text with user-controllable fonts in a zero-shot manner.
@@ 2 Related Work
The majority of algorithms rely on visual glyphs - pixel-level representations that describe the appearance of texts - to guide the generation process. However, massive real-world images in the training dataset often lack ground-truth font annotations, while it is challenging to extract exact font labels from any real-world photos. Consequently, these approaches utilize a fixed standard font to render the texts on their glyph controls. For instance, GlyphControl Yang et al. (2024b) and GlyphDraw Ma et al. (2023) render text detected by OCR models, with the former introducing an additional Glyph ControlNet to regulate the decoder of the original ControlNet model Zhang et al. (2023b). TextMaster Wang et al. (2024a), DiffUTE Chen et al. (2024a), and AnyTrans Qian et al. (2024) aim to maintain uniform fonts within the same image. Choi et al. (2024) explores design translation for artistic characters. Similarly, Zhang et al. (2024) employs randomly chosen fonts and a Canny edge detector. Several approaches also incorporate automatic textual layout generation Tuo et al. (2023); Zhu et al. (2024); Chen et al. (2024a); Li et al. (2024c); Seol et al. (2025); Lakhanpal et al. (2024); Paliwal et al. (2024b); Zhao et al. (2024a). For example, Zhu et al. (2024) uses a vision-language model to identify suitable layouts, DiffUTE Chen et al. (2024a) leverages language models and OCR detectors to propose bounding boxes, and Seol et al. (2025) employs HTML codes. In constrast, we concentrate on human-in-the-loop local text editing rather than automatic global layout generation. Our work also eliminates the need for any large language or multimodal models, streamlining the pipeline.
Our work builds upon the codebase of AnyText Tuo et al. (2023), a glyph-based algorithm that trains a base ControlNet Zhang et al. (2023b) model to render visual texts. Since ground-truth font annotations are not accessible from real-world images, glyphs are generated in a fixed standard font, leaving the model to infer an appropriate font.
@@
Fewer recent works are more closely related to ours in enabling controllable fonts Tuo et al. (2024); Ma et al. (2024); Li et al. (2024a); Shi et al. (2024); Paliwal et al. (2024a); Liu et al. (2025). However, none of these works provide a quantitative evaluation metric to assess the generated fonts in open-world settings. AnyText2 Tuo et al. (2024) is a concurrent work developed by the authors of AnyText Tuo et al. (2023). We share a similar architecture, but unlike Tuo et al. (2024), we eliminate its use of lengthy language prompts in the inputs, separate models to support different languages, and the trainable OCR model to encode the font features. Instead, we use OCR solely to filter out low-quality glyph controls. Tuo et al. (2024) is also not yet open-sourced at the time of our submission.
Glyph-ByT5 Liu et al. (2025) and its v2 Liu et al. (2024) also support font control. However, unlike Liu et al. (2025), which pre-specifies font labels for each language and focuses on improving language understanding, we assume that text rendering is independent of the semantic meanings; in our approach, only pixel-level controls matter. GlyphDraw2 Ma et al. (2024) learns font features through cross-attention between glyph features and hidden variables within the image while fine-tuning a language model to generate font layouts. Unfortunately, it lacks quantitative assessment on fonts, relying on human evaluations. JoyType Li et al. (2024a) focuses on e-commerce product images in non-Latin languages using 10 pre-specified fonts with synthetic images and additional vision-language models. It employs an OCR model for font perceptual losses, while we assume OCR to be font-agnostic. FonTS Shi et al. (2024) uses additional reference images as font style controls and requires users to select fonts from pre-specified integer font labels. Similarly, CustomText Paliwal et al. (2024a) relies on specific font names provided in extended user prompts and an additional training datasets tailored for smaller fonts. While Liu et al. (2024) emphasizes visual aesthetics and highlights challenges with small fonts, our work centers on font editing and share the same perspective that smaller fonts pose greater difficulties. However, we show that zooming into localized regions for text editing can already improve the quality of smaller text. Different from those works, we eliminate the need for pre-specified fonts with unique font names or special font tokens Liu et al. (2024); Shi et al. (2024); Paliwal et al. (2024a) in user prompts, making our method generalizable to unseen languages and fonts.
@@ 3.1 Data-Driven Insights
ControlText employs a data-driven approach to unlock visual text rendering with user-controllable fonts without relying on complex architectural designs, aiming to avoid pitfalls associated with the bitter lesson Sutton (2019) in foundation models. We feed the model a massive and diverse amount of unsupervised glyphs, each containing detailed font features in pixel space, and train the diffusion process to reconstruct target images following the explicit pixel-level hints in these input glyphs.
The key insight is that during training, the model learns to leverage pixel-level controls provided by the glyphs as a shortcut for generating text within images. Importantly, the glyphs no longer need to use a standard font, but can mimic any font appear in the target images. Unlike traditional methods, no annotations of font labels are required; the only guidance comes from how the text appears in pixel space as depicted by the input glyphs.
@@ 3.2.3 Main Training Process
  • Masked image $\boldsymbol{c}_{m}\in\mathbb{R}^{n\times n\times 3}$: An RGB image normalized to the range $[-1,1]$, where the region within the box position $\boldsymbol{c}_{p}$ is masked to 0. Every other pixel is identical to the target image $\boldsymbol{I}\in\mathbb{R}^{n\times n\times 3}$.
@@
$p_{\theta}\left(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{c}\right)=\mathcal{N}\left(\boldsymbol{x}_{t-1};\mu_{\theta}\left(\boldsymbol{x}_{t},t,\boldsymbol{c}\right),\Sigma_{\theta}(t)\right)\vspace{-5mm}$ (2)
@@ 4 Experimental Results
Table 1: Results on AnyText-benchmark. “AnyText-v1.1 Font Aware" refers to AnyText Tuo et al. (2023), but uses the same font-aware glyph controls from our work without fine-tuning as an ablation, which shows significant performance drop in text accuracy. Although with little decrease in non-Latin text accuracy, ControlText outperforms AnyText by a large margin in preserving diverse and detailed font information in input glyph controls.
@@
Methods Text Accuracy $\uparrow$ Image Quality $\downarrow$ Fuzzy Font Accuracy (in Distance) $\downarrow$
[…]
AnyText-v1.1 Tuo et al. (2023) 0.8315 0.9081 26.18 0.3038 0.2581 0.2496 0.2466 0.1615 0.1599 0.1603 0.1614
AnyText-v1.1 Font-Aware 0.5524 0.6839 24.95 0.3049 0.2589 0.2497 0.2458 0.1601 0.1583 0.1587 0.1597
ControlText (ours) 0.8345 0.9107 25.93 0.2449 0.1850 0.1750 0.1709 0.1305 0.1286 0.1291 0.1301
@@
Methods Text Accuracy $\uparrow$ Image Quality $\downarrow$ Fuzzy Font Accuracy (in Distance) $\downarrow$
[…]
AnyText-v1.1 Tuo et al. (2023) 0.8591 0.8284 27.08 0.3103 0.2724 0.2643 0.2608 0.1521 0.1513 0.1519 0.1530
AnyText-v1.1 Font-Aware 0.5578 0.3396 27.11 0.3106 0.2799 0.2713 0.2690 0.1478 0.1472 0.1477 0.1488
ControlText (ours) 0.7867 0.7479 28.63 0.2602 0.1992 0.1891 0.1848 0.1274 0.1271 0.1275 0.1286
@@ 4.1 Experimental Setups
We finetune the ControlNet Zhang et al. (2023b) model, with a size of around $1$B pretrained by AnyText Tuo et al. (2023), for 10 epochs using 4 NVIDIA V100 GPUs, each with 32 GB memory. We use a batch size of 6, a learning rate of $2\times 10^{-5}$, and focus solely on inpainting masked images. The dataset is curated from AnyWord-3M Tuo et al. (2023) but with our font-aware glyph. Each RGB image of size $512$ by $512$ has at most $5$ lines of text. The dataset comprises approximately 1.39 million images in English, 1.6 million images in Chinese, and 10 thousand images in other languages. Following this, we continue training the model for another 2 epochs, turning on the textual perception loss introduced in Tuo et al. (2023). We use AnyText-benchmark Tuo et al. (2023) with 1000 test images in English and Chinese to show quantitative results.
@@ 5 Discussion
A text segmentation model can capture nuanced font information in pixel space without requiring font label annotations in the dataset, enabling zero-shot generation on unseen languages and fonts, as well as scalable training on web-scale image datasets as long as they contain text.
@@
Future work will focus on enhancing data efficiency in the training pipeline, especially fonts in low-resource languages, as well as enabling more complicated artistic style control of text from user prompts beyond font information, including its interaction with diverse underlying background.
@@ 6 Limitations
Additionally, due to limited training resources, our experiments were conducted using a smaller diffusion model as a proof-of-concept compared to commercial ones. Each epoch requires approximately 380 GPU hours on NVIDIA V100 GPUs with 32 GB of memory, but we anticipate significantly improved efficiency on newer hardware and with a larger memory. This constraint may result in suboptimal inpainting of background regions within the text area, as well as instability in the quality of rendered text. The users also have limited controls of background pixels behind the text.

### v2

@@ Abstract
This work demonstrates that diffusion models can achieve font-controllable multilingual text rendering using just raw images without font label annotations. Visual text rendering remains a significant challenge. While recent methods condition diffusion on glyphs, it is impossible to retrieve exact font annotations from large-scale, real-world datasets, which prevents user-specified font control. To address this, we propose a data-driven solution that integrates the conditional diffusion model with a text segmentation model, utilizing segmentation masks to capture and represent fonts in pixel space in a self-supervised manner, thereby eliminating the need for any ground-truth labels and enabling users to customize text rendering with any multilingual font of their choice. The experiment provides a proof of concept of our algorithm in zero-shot text and font editing across diverse fonts and languages, providing valuable insights for the community and industry toward achieving generalized visual text rendering. Code is available at github.com/bowen-upenn/ControlText
@@
Bowen Jiang1 , Yuan Yuan [^1] footnotemark: 1 , Xinyi Bai2, Zhuoqun Hao1, Alyson Yin3, Yaojie Hu1, Wenyu Liao1, Lyle Ungar1, Camillo J. Taylor1 University of Pennsylvania1 Cornell University2 University of California, Irvine3 Philadelphia, PA 19104 Ithaca, NY 14850 Irvine, CA 92697 bwjiang, yyuan86@seas.upenn.edu, ungar@cis.upenn.edu, cjtaylor@seas.upenn.edu
@@ 1 Introduction
Diffusion model is one of the dominant paradigms in image generation Ho et al. (2020); Saharia et al. (2022); Zhang et al. (2023b); Rombach et al. (2022); Betker et al. (2023); Ramesh et al. (2022); Esser et al. (2024), because of its iterative denoising process that allows fine-grained image synthesis. While these models effectively capture data distributions of photorealistic or artistic images, they still fall short in generating high-fidelity text. Rendering text in images is inherently more challenging as it requires precise knowledge of the geometric alignment among strokes, the arrangement of letters as words, the legibility across varying fonts, sizes, and styles, and the integration of text into visual backgrounds. At the same time, humans are more sensitive to minor errors in text, such as a missing character or an incorrectly shaped letter, compared to natural elements in a visual scene that allow for a much higher degree of variation.
@@
  • There are numerous fonts available in the open world, but many fonts with different names are very similar, which confounds evaluation.
@@
We also provide the community with a comprehensive dataset for font-aware glyph controls collected from diverse real-world images. We further propose a quantitative evaluation metric for handling fuzzy fonts in the open world. Experimental results demonstrate that our method, ControlText, as a text and font editing model, facilitates a human-in-the-loop process to generate multilingual text with user-controllable fonts in a zero-shot manner.
@@ 2 Related Work
The majority of algorithms rely on visual glyphs, pixel-level representations of texts, to guide the generation process. However, because most image training datasets lack ground-truth font annotations, most algorithms utilize a fixed standard font to render the texts on their glyph controls. For instance, GlyphControl Yang et al. (2024b) and GlyphDraw Ma et al. (2023) render OCR-detected text using a fixed font, with the former adding a glyph-specific ControlNet Zhang et al. (2023b). TextMaster Wang et al. (2024a), DiffUTE Chen et al. (2024a), and AnyTrans Qian et al. (2024) enforce font consistency within images, while Zhang et al. (2024) introduces font variation through random sampling. Layout generation is also addressed using language and vision models Tuo et al. (2023); Zhu et al. (2024); Li et al. (2024c); Seol et al. (2025); Lakhanpal et al. (2024); Paliwal et al. (2024b); Zhao et al. (2024a). In contrast, we focus on human-in-the-loop text editing without large language or multimodal models.
Our work builds upon the codebase of AnyText Tuo et al. (2023), a glyph-based algorithm that trains a base ControlNet Zhang et al. (2023b) model to render visual texts, with glyphs being generated in a fixed standard font due to unavailability of ground-truth font annotations, leaving the model to infer an appropriate font.
@@
Fewer recent works are more closely related to ours in enabling controllable fonts Tuo et al. (2024); Ma et al. (2024); Li et al. (2024a); Shi et al. (2024); Paliwal et al. (2024a); Liu et al. (2025). However, none of these works provide a quantitative evaluation metric to assess the generated fonts in open-world settings. AnyText2 Tuo et al. (2024) is a concurrent work developed by the authors of AnyText Tuo et al. (2023). We share a similar architecture Tuo et al. (2024), but we eliminate its use of lengthy language prompts in the inputs, separate models to support different languages, and the trainable OCR model to encode the font features. Instead, we use OCR solely to filter out low-quality glyph controls. Tuo et al. (2024) is also not yet open-sourced at the time of our submission.
Several works tackle font control through predefined font labels or additional supervision. Glyph-ByT5 Liu et al. (2025, 2024) requires language-specific font labels and emphasizes language understanding, while we treat text rendering as purely visual. GlyphDraw2 Ma et al. (2024) learns font features via cross-attention and a fine-tuned language model, but lacks quantitative font evaluation. JoyType Li et al. (2024a) focuses on synthetic e-commerce images with 10 fixed fonts and vision-language models, assuming OCR is font-sensitive—unlike our font-agnostic assumption. FonTS Shi et al. (2024) and CustomText Paliwal et al. (2024a) rely on pre-specified font labels or user-defined font names. While Liu et al. (2024) highlights challenges with small fonts, we show that localized editing improves small-font quality. In contrast to above methods, we eliminate the need for font labels, special tokens, or predefined font names Liu et al. (2024); Shi et al. (2024); Paliwal et al. (2024a), enabling generalization to unseen fonts and languages.
@@ 3.1 Data-Driven Insights
ControlText enables user-controllable font rendering through a simple, data-driven approach, without complex architectures, embracing the principles of the bitter lesson Sutton (2019). By training on diverse unsupervised glyphs rich in pixel-level font details, the diffusion model learns to reconstruct images directly from visual cues. The key insight is that the model learns to use pixel-level glyphs as direct cues for text generation, eliminating the need for font labels. Glyphs can mimic any target font, with guidance provided solely by their visual appearance.
@@ 3.2.3 Main Training Process
  • Masked image $\boldsymbol{c}_{m}\in\mathbb{R}^{n\times n\times 3}$: An RGB image normalized to the range [1,1][-1,1], where the region within the box position $\boldsymbol{c}_{p}$ is masked to 0. Every other pixel is identical to the target image $\boldsymbol{I}\in\mathbb{R}^{n\times n\times 3}$.
@@
$p_{\theta}\left(\boldsymbol{x}_{t-1}\mid\boldsymbol{x}_{t},\boldsymbol{c}\right)=\mathcal{N}\left(\boldsymbol{x}_{t-1};\mu_{\theta}\left(\boldsymbol{x}_{t},t,\boldsymbol{c}\right),\Sigma_{\theta}(t)\right)\vskip-14.22636pt$ (2)
@@ 4 Experimental Results
Table 1: Results on AnyText–benchmark. “AnyText with font–aware glyphs” denotes AnyText Tuo et al. (2023) that directly adopts our glyph controls without fine-tuning as an ablation. ControlText consistently preserves detailed font information while maintaining strong text accuracy.
@@
Method Text Accuracy $\uparrow$ Fuzzy Font Accuracy (distance) $\downarrow$
SenACC NED $l_{2}@5$ $l_{2}@20$ $l_{2}@50$ $l_{2}@\text{full}$ $\cos@5 cos$$\cos@20$ $\cos@50 cos$$\cos@{\text{full}}$
ControlText (ours) 0.8345 0.9537 0.3431 0.3387 0.3382 0.3381 0.2710 0.2523 0.2504 0.2500
AnyText Tuo et al. (2023) 0.8315 0.9518 0.4654 0.4628 0.4623 0.4622 0.4125 0.3974 0.3954 0.3948
AnyText w/ glyphs 0.5524 0.8387 0.4738 0.4709 0.4705 0.4703 0.4261 0.4096 0.4077 0.4070
TextDiffuser Chen et al. (2024b) 0.5966 0.8236 0.6293 0.6280 0.6278 0.6277 0.5536 0.5448 0.5436 0.5432
GlyphControl Yang et al. (2023) 0.4098 0.7414 0.6046 0.6031 0.6029 0.6028 0.5423 0.5321 0.5307 0.5302
ControlNet Zhang et al. (2023b) 0.5837 0.8305 0.6853 0.6839 0.6837 0.6837 0.5896 0.5812 0.5802 0.5798
@@
Method Text Accuracy $\uparrow$ Fuzzy Font Accuracy (distance) $\downarrow$
SenACC NED $l_{2}@5$ $l_{2}@20$ $l_{2}@50$ $l_{2}@\text{full}$ $\cos@5 cos$$\cos@20$ $\cos@50 cos$$\cos@{\text{full}}$
ControlText (ours) 0.7867 0.9276 0.3561 0.3508 0.3499 0.3497 0.3295 0.3015 0.2975 0.2964
AnyText Tuo et al. (2023) 0.8591 0.9515 0.4632 0.4593 0.4585 0.4583 0.4743 0.4488 0.4444 0.4431
AnyText w/ glyphs 0.5578 0.8120 0.4683 0.4646 0.4638 0.4636 0.4808 0.4553 0.4512 0.4490
TextDiffuser Chen et al. (2024b) 0.0611 0.2816 0.6773 0.6761 0.6757 0.6756 0.6638 0.6528 0.6509 0.6502
GlyphControl Yang et al. (2023) 0.0377 0.2338 0.7298 0.7285 0.7282 0.7281 0.6995 0.6891 0.6873 0.6865
ControlNet Zhang et al. (2023b) 0.3500 0.6393 0.7609 0.7594 0.7591 0.7590 0.7123 0.7018 0.7001 0.6994
Table 2: Fuzzy font accuracy on held-out Kannada and Korean languages in zero-shot settings. Lower values indicate better font similarity preservation. ControlText consistently outperforms AnyText with better generalization to completedly unseen languages, without explicitly being trained on them.
Method Fuzzy Font Accuracy (distance) $\downarrow$
$l_{2}@5$ $l_{2}@20$ $l_{2}@50$ $l_{2}@\text{full}$ $\cos@5 cos$$\cos@20$ $\cos@50 cos$$\cos@{\text{full}}$
ControlText (ours) 0.4564 0.6090 0.4507 0.6075 0.4503 0.6074 0.4503 0.6073
AnyText Tuo et al. (2023) 0.4587 0.6209 0.4526 0.6193 0.4522 0.6192 0.4521 0.6192
@@ 4.1 Experimental Setups
We finetune the ControlNet Zhang et al. (2023b) model, with a size of around 11B pretrained by AnyText Tuo et al. (2023) and several convolutional encoders to process input conditions, for 10 epochs using 4 NVIDIA V100 GPUs, each with 32 GB memory. We use a batch size of 6, a learning rate of $2\times 10^{-5}$, and focus solely on inpainting masked images. The dataset is curated from AnyWord-3M Tuo et al. (2023) but with our font-aware glyph. Each RGB image of size $512$ by $512$ has at most $5$ lines of text. The dataset comprises approximately 1.39 million images in English, 1.6 million images in Chinese, and 10 thousand images in other languages. Following this, we continue training the model for another 2 epochs, turning on the textual perception loss introduced in Tuo et al. (2023). We use AnyText-benchmark Tuo et al. (2023) with 1000 test images in English and Chinese to show quantitative results.
@@ 4.3 Quantitative Results
To further evaluate the cross-lingual generalization ability of our model, we conducted zero-shot inferences on Kannada and Korean scripts from the MLe2e dataset Gomez and Karatzas (2016), previously unseen by the model. As shown in Table 2, ControlText consistently outperforms AnyText across all metrics. Specifically, ControlText achieves lower average cosine and l2 distances across top-5, top-20, top-50, and full-set comparisons in font accuracy metrics, indicating better font-style consistency and generation quality. Meanwhile, these results reinforce that ControlText generalizes better to underrepresented scripts such as Kannada and Korean, even without explicit training on these languages.
@@ 5 Discussion
A text segmentation model can capture nuanced font information in pixel space without requiring font label annotations in the dataset, enabling zero-shot generation on unseen languages and fonts (as evidenced by generated images with Japanese in Figure 1), as well as scalable training on web-scale image datasets as long as they contain text.
@@
Future work will focus on improving data efficiency in the training pipeline, particularly for fonts in low-resource languages. We also plan to explore reinforcement learning with text and font verification signals in multimodal transformers. Additionally, we aim to enable more advanced artistic style control of text based on user prompts, extending beyond font attributes to include interactions with diverse background content.
@@ 6 Limitations
Additionally, due to limited training resources, our experiments were conducted using a smaller diffusion model as a proof-of-concept compared to commercial ones. Each epoch requires approximately 380 GPU hours on NVIDIA V100 GPUs with 32 GB of memory, but we anticipate significantly improved efficiency on newer hardware and with a larger memory. This constraint may result in suboptimal inpainting of background regions within the text area, as well as instability in the quality of rendered text. This also made it difficult to try other base models for comparison. Moreover, the limited training resources also made it difficult to conduct more ablation studies, which could have provided more nuanced insights into the architecture’s effectiveness. The users also have limited controls of background pixels behind the text.
@@ 7 Ethical Impact
8 Acknowledgment
This work was supported by the National Science Foundation (NSF) grant CCF-2112665 (TILOS).